Suppose a firm’s tax rate is 35%. a. What effect would a $7 million operating expense have on this year’s earnings? What effect would it have on next year’s earnings? b. What effect would a $7 million capital expense have on this year’s earnings, if the capital is depreciated straightline over 5 years? What effect would it have on next year’s earnings? (Below are all figures in thousand dollar. Round to nearest thousand. If number is negativ use - . Don't use , )

Year 1 Year 2
a. One-time expense -7000
Effect on earnings
b. Capitalized expense -7000
Annual depreciation
Effect on earnings
Answer #1

Q. a). i). Effect on earnings in Year 1 = One time expense - Tax rate * One time expense.

= 7 Million - 35 % of 7 Million.

= 7 Million - 2.45 Million.

= $ 4.55 Million.

ii). Effect on earnings in Year 2 = No effects.

Conclusion :- a). If expense of $ 7 million considered as one time expense only :-

a). i). Effect on earnings in Year 1 $ 4.55 Million.
a). ii). Effect on earnings in Year 2 No effects.

Q. b). i). Effect on earnings in Year 1 = Tax rate * Total expense amount / Number of years depreciated.

= 0.35 * 7 Million / 5

= 0.35 * 1.40 Million

= $ 0.49 Million.

ii). Effect on next year earnings = Tax rate * Total expense amount / Number of years depreciated.

= 0.35 * 7 Million / 5

= 0.35 * 1.40 Million

= $ 0.49 Million.

Conclusion :- b). If expense of $ 7 million considered as capitalized expense only :-

a). i). Effect on earnings in Year 1 $ 0.49 Million.
a). ii). Effect on earnings in Year 2 $ 0.49 Million.